OpenStudy (anonymous):

If h=x^2-1/x+1 + x^2-1/x-1 what is x

hartnn (hartnn):

do u know how to factor x^2-1 ??

OpenStudy (anonymous):

Dont you change it to x^2-1^2

hartnn (hartnn):

yes, and apply a^2-b^2 = (a+b)(a-b) what do u get here ?

OpenStudy (anonymous):

(x+1)(x-1)/(x+1) right?

hartnn (hartnn):

right ! good. so what cancels out ?

OpenStudy (anonymous):

both of the x+1 so your left with x-1

hartnn (hartnn):

thats correct :) and what remains in 2nd term ?

OpenStudy (anonymous):

x+1

hartnn (hartnn):

now add x-1+x+1 = ?

OpenStudy (anonymous):

is it just 2x

hartnn (hartnn):

so h=2x hence x= h/2 got it ?

OpenStudy (anonymous):

yes thank you

hartnn (hartnn):

welcome :)